OpenStudy (anonymous):

Given that f(x)=(3x^2)-5, how do I find (f(x+h)-f(x))/h and (f(w)-f(x))/(w-x)?

OpenStudy (helder_edwin):

for example. since u have \[ \large f(x)=3x^2-5 \] then \[ \large f(A)=3A^2-5 \] \[ \large f(a+b)=3(a+b)^2-5 \]

OpenStudy (anonymous):

Thanks!